# Towards the Worst-case Robustness of Large Language Models

### v3

@@ Appendix D Proofs For Value-to-weight Ratio and Volume for Specific Kernels
The first way to express this combination number is by first considering the tokens that differ from the corresponding part of $\bm{x}_{adv}$ but match $\bm{x}$. These tokens account for $\binom{d}{d-i}$. Among the remaining $i$ tokens, $i+j-d$ tokens must differ from both $\bm{x}_{adv}$ and $\bm{x}$, so they contribute $\binom{i}{i+j-d}$. The remaining tokens differ from the corresponding part of $\bm{x}$ but match $\bm{x}_{adv}$. Therefore, we have:
@@
Similarly, we can express this combination number from the perspective of $\bm{x}_{adv}$ instead of $\bm{x}$. First, we consider the tokens that differ from the corresponding part of $\bm{x}$ but match $\bm{x}_{adv}$. These tokens contribute $\binom{d}{d-j}$. Among the remaining $j$ tokens, $i+j-d$ tokens must differ from both $\bm{x}_{adv}$ and $\bm{x}$, contributing $\binom{j}{i+j-d}$. The remaining tokens differ from the corresponding part of $\bm{x}_{adv}$ but match $\bm{x}$. Thus, we get:
@@
These two combinations are actually the same, as shown by the symmetrization lemma in Theorem E.1. This symmetry provides many favorable properties for the uniform kernel.
@@ D.1.2 Case Study: $d=2$
When $d=2$, there are $93^{2}=$ 9 cases.
@@ D.1.3 Case study: $d=3$
This simple result enables us to intuitively illustrate the greedy algorithm using $p_{adv}-p_{A}$ graph. See Sec. D.3 and Figure 1(a) for detail.
One can also interpret the certified bound for absorbing kernel in another way, similar to [29]:
@@
Therefore, the output $g(\bm{x}_{adv})$ must satisfy $g(\bm{x}_{adv})\geq\sum_{\bm{z}\in L_{1}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$. Note that for $\bm{z}\in L_{1}$, $p(\bm{z}|\bm{x}_{adv})=p(\bm{z}|\bm{x})$, so Theorem 5.2 holds. Additionally, there exists a worst-case $f$ where $f=0$ for all $\bm{z}\in L_{2}$, making this bound tight.
@@ D.3 Analytic Solution of Certified Robustness using Absorbing Kernel
We analyze the $p_{adv}-p_{A}$ plots (where $p_{adv}$ is on the vertical axis and $p_{A}$ is on the horizontal axis), which provide a direct illustration of the Knapsack algorithm. As shown in Figure 1(a), $p_{adv}=0$ when $p_{A}\leq 1-\beta^{d}$. When $p_{A}\geq 1-\beta^{d}$, we trade $p_{adv}$ for $p_{A}$ at a trading rate of 1 (indicated by a slope of 1).
@@
D.4 Proof of Theorem 5.5
Since the certified robustness of the uniform kernel does not have an analytic solution, proving Theorem 5.5 requires some subtle observations.
@@
The condition $\left(\frac{\alpha}{\bar{\beta}}\right)\leq 1$ is equivalent to $\bar{\beta}\geq\frac{1}{\mathcal{V}}$, and this is always satisfied because at $t_{\max}$ the uniform prior assigns equal probability $\frac{1}{\mathcal{V}}$ to all tokens. Therefore, Eq. 10 provides a sufficient condition for Theorem 5.5.
In the following subsections, we first present a complete proof of Eq. 10. Then, we analyze some simple cases to provide intuition on how we arrive at this proof.
D.4.1 Final Proof of Sufficient Condition Eq. 10
@@
Using this lemma, we upper bound Eq. 10 by:
@@
Which completes the proof of Eq. 10. Since Eq. 10 is a sufficient condition of Theorem 5.5, this also completes the proof of Theorem 5.5.
@@
The key insight here is that $\alpha^{i}=\frac{\beta^{i}}{(|\mathcal{V}|-1)^{i}}$, contain a high order term $\frac{1}{(|\mathcal{V}|-1)^{i}}$. We know that $i+j-d\leq i$ since $j\leq d$. When $i+j-d<i$, $(|\mathcal{V}|-2)^{i+j-d}\alpha^{i}=(|\mathcal{V}|-2)^{i+j-d}\frac{\beta^{i}}{(|\mathcal{V}|-1)^{i}}\to 0$. Hence, we only need to consider $i+j-d=i$, or equivalently, $j=d$. Therefore, we have the following:
@@
Intuitively, the certified robustness would be the smallest when $|\mathcal{V}|\to\infty$. This inspired us to bound Eq. 10 using $j=d$. However, the last step $\frac{(|\mathcal{V}|-2)^{d}}{(|\mathcal{V}|-1)^{d}}=1$ does not hold when $|\mathcal{V}|\neq\infty$. Therefore, we consider loosing by $i=d$ when proving Eq. 10. This case study also demonstrates that Eq. 10 is almost tight since it becomes equality when $|\mathcal{V}|\to\infty$.
When $|\mathcal{V}|\to\infty$, the value-to-weight ratio $\frac{\alpha^{i}\bar{\beta}^{d-i}}{\alpha^{j}\bar{\beta}^{d-j}}=\frac{\alpha^{i-j}}{\bar{\beta}^{i-j}}=\left(\frac{\alpha}{\bar{\beta}}\right)^{i-j}$ only have three possible values: 0 when $i>j$, 1 when $i=j$, $\infty$ when $i<j$. Since for all $p_{A}\leq 1-\beta^{d}$, we have $i>j$, thus $p_{adv}=0$ for all $p_{A}\leq 1-\beta^{d}$. By symmetrization lemma (Theorem E.1), $i=j$ must hold for all $p_{A}\geq 1-\beta^{d}$. Therefore, the $p_{adv}-p_{A}$ graph of the uniform kernel and absorbing kernel is exactly the same. This means Figure 1(b) gradually goes to Figure 1(a) when $|\mathcal{V}|\to\infty$.
D.4.3 Simple case study: $d$=1,2,3
@@
$\sum_{0\leq i<j\leq d}v(i,j)=v(0,3)+v(1,3)+v(1,2)+v(2,3)=\bar{\beta}^{3}+3(|\mathcal{V}|-2)\bar{\beta}^{2}\alpha+3\bar{\beta}^{2}\alpha+3(|\mathcal{V}|-2)^{2}\bar{\beta}\alpha^{2}$

### v4

@@ Appendix D Proofs For Value-to-weight Ratio and Volume for Specific Kernels
The first way to express this combination number is by first considering the tokens that differ from the corresponding part of $\bm{x}_{adv}$ but match $\bm{x}$. These tokens account for $\binom{d}{d-i}$. Among the remaining $i$ tokens, i+jdi+j-d tokens must differ from both $\bm{x}_{adv}$ and $\bm{x}$, so they contribute $\binom{i}{i+j-d}$. The remaining tokens differ from the corresponding part of $\bm{x}$ but match $\bm{x}_{adv}$. Therefore, we have:
@@
Similarly, we can express this combination number from the perspective of $\bm{x}_{adv}$ instead of $\bm{x}$. First, we consider the tokens that differ from the corresponding part of $\bm{x}$ but match $\bm{x}_{adv}$. These tokens contribute $\binom{d}{d-j}$. Among the remaining $j$ tokens, i+jdi+j-d tokens must differ from both $\bm{x}_{adv}$ and $\bm{x}$, contributing $\binom{j}{i+j-d}$. The remaining tokens differ from the corresponding part of $\bm{x}_{adv}$ but match $\bm{x}$. Thus, we get:
@@
These two combinations are actually the same, as shown by the symmetrization lemma in Theorem˜E.1. This symmetry provides many favorable properties for the uniform kernel.
@@ D.1.2 Case Study: $d=2$
When $d=2$, there are $3^{2}=9$ cases.
@@ D.1.3 Case study: $d=3$
This simple result enables us to intuitively illustrate the greedy algorithm using $p_{adv}-p_{A}$ graph. See Sec.˜D.3 and Figure˜1(a) for detail.
One can also interpret the certified bound for absorbing kernel in another way, similar to Zeng et al. (2023):
@@
Therefore, the output $g(\bm{x}_{adv})$ must satisfy $g(\bm{x}_{adv})\geq\sum_{\bm{z}\in L_{1}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$. Note that for $\bm{z}\in L_{1}$, $p(\bm{z}|\bm{x}_{adv})=p(\bm{z}|\bm{x})$, so Theorem˜5.2 holds. Additionally, there exists a worst-case $f$ where $f=0$ for all $\bm{z}\in L_{2}$, making this bound tight.
@@ D.3 Analytic Solution of Certified Robustness using Absorbing Kernel
We analyze the $p_{adv}-p_{A}$ plots (where $p_{adv}$ is on the vertical axis and $p_{A}$ is on the horizontal axis), which provide a direct illustration of the Knapsack algorithm. As shown in Figure˜1(a), $p_{adv}=0$ when $p_{A}\leq 1-\beta^{d}$. When $p_{A}\geq 1-\beta^{d}$, we trade $p_{adv}$ for $p_{A}$ at a trading rate of 1 (indicated by a slope of 1).
@@
D.4 Proof of Theorem˜5.5
Since the certified robustness of the uniform kernel does not have an analytic solution, proving Theorem˜5.5 requires some subtle observations.
@@
The condition $\left(\frac{\alpha}{\bar{\beta}}\right)\leq 1$ is equivalent to $\bar{\beta}\geq\frac{1}{\mathcal{V}}$, and this is always satisfied because at $t_{\max}$ the uniform prior assigns equal probability $\frac{1}{\mathcal{V}}$ to all tokens. Therefore, Eq.˜11 provides a sufficient condition for Theorem˜5.5.
In the following subsections, we first present a complete proof of Eq.˜11. Then, we analyze some simple cases to provide intuition on how we arrive at this proof.
D.4.1 Final Proof of Sufficient Condition Eq.˜11
@@
Using this lemma, we upper bound Eq.˜11 by:
@@
Which completes the proof of Eq.˜11. Since Eq.˜11 is a sufficient condition of Theorem˜5.5, this also completes the proof of Theorem˜5.5.
@@
The key insight here is that $\alpha^{i}=\frac{\beta^{i}}{(|\mathcal{V}|-1)^{i}}$, contain a high order term $\frac{1}{(|\mathcal{V}|-1)^{i}}$. We know that $i+j-d\leq i$ since $j\leq d$. When i+jd<ii+j-d<i, $(|\mathcal{V}|-2)^{i+j-d}\alpha^{i}=(|\mathcal{V}|-2)^{i+j-d}\frac{\beta^{i}}{(|\mathcal{V}|-1)^{i}}\to 0$. Hence, we only need to consider i+jd=ii+j-d=i, or equivalently, $j=d$. Therefore, we have the following:
@@
Intuitively, the certified robustness would be the smallest when $|\mathcal{V}|\to\infty$. This inspired us to bound Eq.˜11 using $j=d$. However, the last step $\frac{(|\mathcal{V}|-2)^{d}}{(|\mathcal{V}|-1)^{d}}=1$ does not hold when $|\mathcal{V}|\neq\infty$. Therefore, we consider loosing by $i=d$ when proving Eq.˜11. This case study also demonstrates that Eq.˜11 is almost tight since it becomes equality when $|\mathcal{V}|\to\infty$.
When $|\mathcal{V}|\to\infty$, the value-to-weight ratio $\frac{\alpha^{i}\bar{\beta}^{d-i}}{\alpha^{j}\bar{\beta}^{d-j}}=\frac{\alpha^{i-j}}{\bar{\beta}^{i-j}}=\left(\frac{\alpha}{\bar{\beta}}\right)^{i-j}$ only have three possible values: 0 when $i>j$, 1 when $i=j$, $\infty$ when $i<j$. Since for all $p_{A}\leq 1-\beta^{d}$, we have $i>j$, thus $p_{adv}=0$ for all $p_{A}\leq 1-\beta^{d}$. By symmetrization lemma (Theorem˜E.1), $i=j$ must hold for all $p_{A}\geq 1-\beta^{d}$. Therefore, the $p_{adv}-p_{A}$ graph of the uniform kernel and absorbing kernel is exactly the same. This means Figure˜1(b) gradually goes to Figure˜1(a) when $|\mathcal{V}|\to\infty$.
D.4.3 Simple case study: dd=1,2,3
@@
$\sum_{0\leq i<j\leq d}v(i,j)=v(0,3)+v(1,3)+v(1,2)+v(2,3)$
$=$ $\bar{\beta}^{3}+3(|\mathcal{V}|-2)\bar{\beta}^{2}\alpha+3\bar{\beta}^{2}\alpha+3(|\mathcal{V}|-2)^{2}\bar{\beta}\alpha^{2}$